OpenStudy (anonymous):

What is the first step in solving 3x - 19 = 28? A. Divide by 3 on both sides. B. Multiply by 3 on both sides. C. Add 19 on both sides. D. Subtract 28 on both sides.

OpenStudy (anonymous):

first you try to get all the x's on one side and all the numbers on the other

OpenStudy (anonymous):

So D ?

OpenStudy (mathstudent55):

Look at the variable you are lookig for. It's x. If you had a number for x and you wanted to know what the left sides is equal to you would 1. multiply x by 3 2. subtract 19 In order to solve for x, do the opposite of the operations above in the opposite order. So start by 1. add 19 2. divide by 3

OpenStudy (anonymous):

if you don't understand that ^, I can explain it to you.

OpenStudy (anonymous):

Here, 3x - 19 = 28 take , -19 to the right hand side.. and change it's sign.. 3x=28+19 3x=47 x=47/3 ..